Given f(2) = 3, g(3) = 2, f(3) = 4 and g(2) = 5, evaluate
f + g, f/g

Answer by Edwin McCravy(20081)

Given f(2) = 3, g(3) = 2, f(3) = 4 and g(2) = 5, evaluate
f + g, f/g

(f+g)(2) = f(2)+g(2) = 3+5 = 8

(f+g)(3) = f(3)+g(3) = 4+2 = 6

(f/g)(2) = f(2)/g(2) = 3/5

(f/g)(3) = f(3)/g(3) = 4/2 = 2

Edwin